OpenStudy (anonymous):

the length of the pojection of a line on a plane is 12 cm. if the length of the line is 18cm, find the angle it makes with the plane

OpenStudy (raden):



OpenStudy (raden):

then use the defenition of cosine : cosθ = 12/18 cosθ = 2/3 solve for θ by take the arc(cos(2/3)) (use the calculator)

OpenStudy (anonymous):

thanks a lot :)

OpenStudy (raden):

yo;re welcome